# Action-OOD: An End-to-End Skeleton-Based Model for Robust Out-of-Distribution Human Action Detection

### v1

@@ 4.3.3 Ablation study
Effect of Energy Loss: Comparing the results with only traditional cross-entropy loss (bottom two rows), it can be observed that the method utilizing energy loss employed in this paper significantly enhances the model’s capabilities across various settings. Additionally, we discover that compared with directly employing ASH, the model trained with the feature fusion module but using only cross-entropy loss performs less effectively on OOD tasks, yet slightly outperforms in ID classification. This further indicates that the combined usage of the feature fusion module and energy loss can address the overall capabilities of the model, leading to better differentiation between in-distribution and out-of-distribution data.
@@ 4.3.4 Different pruning percentages
To explore the effect of pruning percentage $p$, different values are tested in our experiments. As shown in Fig. 5, we select the best-performing ASH-P strategy on the NTU60 data set for evaluation. The size of the blue points reflects the size of the pruning ratio $p$. The red, green, and purple points respectively represent the performance of the ReAct, DICE, and DICE+ReAct methods at their optimal ratios.
Figure 5: ID-OOD tradeoff on NTU-RGB+D 60 dataset. The size of the points in the enlarged line chart corresponds to different $p$ values. The line chart represents the results of our method taking p-values in the range of 40%-90%.
@@ 5 Conclusion and future works
This paper introduces Action-OOD, an end-to-end framework designed to address the challenge of recognizing out-of-distribution human action samples. The framework adopts a hierarchical graph composition method and utilizes GCN for extracting the feature maps of human skeletons. Subsequently, a feature activation strategy is applied and we fuse the activated features with the original ones through a proposed feature fusion block. Experiments show that this step helps to equip the model with the innate ability to identify out-of-distribution data while still maintaining the capability of ID classification. Next, an energy score value is computed for each test sample based on the logits output by the classifier. This helps determine whether it belongs to an OOD category. The model is trained using a designed energy-based loss function, which helps the model to reduce the variance of ID data and differentiate ID and OOD samples. Experimental results demonstrate that this approach effectively mitigates the overconfidence issue encountered by human action classification models when dealing with samples beyond the training category. Notably, since the model training process relies exclusively on sample information within the branch without prior knowledge of OOD distribution characteristics, it exhibits broad practical applicability.
@@
Admittedly, the model still needs to address some potential limitations. Firstly, since the model is trained on GCN, its generalization ability is constrained by the number of skeleton nodes across various datasets. In real-world scenarios, if the trained model is deployed to identify whether a test sample with a different number of nodes is out-of-distribution, its accuracy may be compromised. Secondly, the end-to-end model’s generalization capability is still bounded by the model parameters and the size of the dataset. One potential solution is to explore semi-supervised learning techniques to mitigate this issue. Moreover, the problem can be further addressed by integrating a large language model and conducting training on datasets with limited ID samples. By leveraging existing category text and feature information, the model can infer the name of the unknown category for unfamiliar sample features. In summary, human action recognition tasks present diverse application scenarios, with numerous novel challenges awaiting exploration based on specific practical contexts.

### v2

@@ 4.3.2 Number of dimensions for OOD representation
4.3.3 Other loss functions for OOD detection
To evaluate the effectiveness of the proposed energy-based loss function, we compare it with another method LogitNorm [41]. Experiments use our fixed framework but training with the two different loss functions. Results for both OOD detection and ID recognition tasks are shown in Table 5 and Table 6, respectively.
Table 5: Results on ID-OOD mixed dataset trained with LogitNorm loss funcion. The best results of each column are in bold.
OOD NTU60 NTU120 Kinetics400 Average
Methods
Error ↓
FPR95 ↓
AUROC ↑
Error ↓
FPR95 ↓
AUROC ↑
Error ↓
FPR95 ↓
AUROC ↑
Error ↓
FPR95 ↓
AUROC ↑
LogitNorm
ASH-P
34.33 63.66 78.22 30.75 56.50 84.57 51.23 88.73 52.69 38.77 69.63 71.83
LogitNorm
ASH-S
39.11 73.21 76.50 41.09 77.19 75.01 51.39 96.74 46.55 43.86 82.38 66.02
LogitNorm
ASH-B
43.10 81.21 75.05 32.85 60.71 86.70 51.46 94.69 52.70 51.09 96.10 51.31
Table 6: Results on ID sample-only test dataset with ST-GCN extractor. The best results of each column are in bold.
ID sample-only NTU60 NTU120 Kinetics400 Average
Methods Top1 ↑ Overlap ↓ Top1 ↑ Overlap ↓ Top1 ↑ Overlap ↓ Top1 ↑ Overlap ↓
LogitNorm-ASH-P 90.66 0.37 89.05 0.47 26.30 0.92 68.67 0.59
LogitNorm-ASH-S 90.39 0.59 88.17 0.64 21.39 0.93 66.76 0.65
LogitNorm-ASH-B 4.44 0.79 17.63 0.79 26.19 0.91 16.09 0.83
By comparison, it can be seen that in most experiments, the model trained with our proposed energy-based loss function outperforms the results obtained with LogitNorm. It can also be observed that the results trained with LogitNorm show significant variation among different activation methods, particularly with the ASH-B method. We speculate that this may be because the LogitNorm method itself is a correction approach based on the original softmax prediction activation. Moreover, our designed framework partly uses ASH to activate the skeleton features and then applies the LogitNorm loss function for correction. This may lead to either insignificant effects (as seen in LogitNorm-ASH-P and LogitNorm-ASH-S) or overcorrection (as seen in LogitNorm-ASH-B). On the other hand, our framework utilizes ASH for feature activation, and during the training stage, the proposed loss function directly reinforces the learning of the OOD discrimination target through energy distance. The experimental results also show that for our framework, using this target-enhancing training method during the training stage through loss function is more effective. Therefore, overall, our framework not only addresses the issue of OOD feature extraction from the perspective of feature activation but also strengthens the range of ID sample energy scores from the perspective of the energy value distribution with the aim of discrimination. This multi-angle approach enables the model to address both OOD and ID sample recognition problems effectively.
@@ 4.3.4 Ablation study
  • 3.
    Effect of Energy Loss: Comparing the results with only traditional cross-entropy loss (bottom two rows), it can be observed that the method utilizing energy loss employed in this paper significantly enhances the model’s capabilities across various settings. Additionally, we discover that compared with directly employing ASH, the model trained with the feature fusion module but using only cross-entropy loss performs less effectively on OOD tasks, but slightly outperforms in ID classification. This further indicates that the combined use of the feature fusion module and energy loss can address the overall capabilities of the model, leading to better differentiation between in-distribution and out-of-distribution data.
Figure 5: Distributions of ID and OOD samples among different ablation studies. The blue area indicates the ID distribution, and the orange shows the OOD samples. The value of overlap represents the amount of overlapping area between the two distributions.
  • 1.
    The impact of different model components on energy score distribution: From Figure 7, it can be observed that incorporating the ASH method during the training stage (Figure 7(b)) reduces the range of energy score distribution compared to the pure GCN model. Additionally, it creates the distance between the two distribution peaks. Contrasting Figure 7(b) with Figure 7(d), where the only difference lies in the usage of a feature fusion block, shows a further reduction in the distribution range from 20 to 12. Moreover, the variance of the ID data is also diminished. Furthermore, as evident from Figure 7(c) and (d), adjustments to the Energy-based loss function result in not only a decreased range of the data but also a reduction in the variance of the OOD data. The combination of these effects minimizes the overlap area of the ID and OOD distributions in our model test results.
@@ 4.3.5 Different pruning percentages
Figure 6: ID-OOD tradeoff on the NTU-RGB+D 60 dataset. The size of the points in the enlarged line chart corresponds to different $p$ values. The line chart represents the results of our method for p-values in the range of 40%-90%.
To explore the effect of pruning percentage $p$, different values are tested in our experiments. As shown in Figure 6, we select the best-performing ASH-P strategy on the NTU60 dataset for evaluation. The size of the blue points reflects the pruning ratio $p$. The red, green, and purple points respectively represent the performance of the ReAct, DICE, and DICE+ReAct methods at their optimal ratios.
@@
4.3.6 Different backbones of skeleton feature extractor
As stated in section 1, we focus on exploiting the extracted features to extend their adaptability for OOD detection while maintaining robust prediction performance for the supervised classes. Thus, we use the state-of-the-art supervised skeleton-based recognition model HD-GCN as feature extractor. The results demonstrate that the model has achieved strong OOD recognition performance compared to many other baselines.
Additionally, we selected the ST-GCN model, which has a similar framework but underperforms compared to HD-GCN in classic action recognition tasks, to serve as a feature extractor for comparative experiments.
Table 8: Results on ID-OOD mixed dataset with ST-GCN extractor. The best results of each column are in bold.
OOD NTU60 NTU120 Kinetics400 Average
Methods
Error ↓
FPR95 ↓
AUROC ↑
Error ↓
FPR95 ↓
AUROC ↑
Error ↓
FPR95 ↓
AUROC ↑
Error ↓
FPR95 ↓
AUROC ↑
ST-GCN
only
35.00 65.01 81.16 56.07 68.36 53.85 51.37 95.11 51.79 47.48 76.16 62.27
ST-GCN
ASH-P
32.47 59.95 84.45 51.68 71.49 71.93 51.90 92.73 56.89 45.35 74.72 71.09
ST-GCN
ASH-S
36.83 68.67 76.63 32.96 60.93 82.90 51.82 93.09 54.85 40.54 74.23 71.46
ST-GCN
ASH-B
43.10 81.21 75.05 32.85 60.71 86.70 51.92 92.67 57.08 42.62 78.20 72.94
Table 9: Results on ID sample-only test dataset with ST-GCN extractor. The best results of each column are in bold.
ID sample-only NTU60 NTU120 Kinetics400 Average
Methods Top1 ↑ Overlap ↓ Top1 ↑ Overlap ↓ Top1 ↑ Overlap ↓ Top1 ↑ Overlap ↓
ST-GCN only 90.33 0.53 48.11 0.81 24.03 0.92 54.16 0.75
ST-GCN-ASH-P 90.97 0.43 88.99 0.72 21.93 0.89 67.30 0.68
ST-GCN-ASH-S 90.08 0.56 88.81 0.46 21.39 0.93 66.76 0.65
ST-GCN-ASH-B 90.41 0.61 87.97 0.41 21.54 0.89 66.64 0.64
Compared with the results of Table 2 and Table 8, on the OOD detection task, the performance of ST-GCN as the backbone feature extractor is not as good as that of Skeleton-OOD. On some datasets, even the best results from ST-GCN do not surpass the worst results from Skeleton-OOD. Although replacing the feature extractor with a less effective one degrades OOD recognition, the results still generally outperform those of other baselines. The results are similar in the ID task (Table 9): ST-GCN’s performance is generally lower than that of Skeleton-OOD and most baseline methods using HD-GCN as the prior skeleton feature extractor. This aligns with findings from numerous existing studies in the field of action recognition, which show that ST-GCN is inferior to HD-GCN. These indicate that ST-GCN, as a backbone, does not perform as well as Skeleton-OOD in most indicators. This further confirms that a superior feature extractor significantly aids in improving OOD recognition.
Also, comparing the results of only ST-GCN and the below activation-added methods in Table 8 and Table 9, it can be concluded that adding activation operations after the feature extractor can indeed improve the OOD detection capability of the model. This is not only limited to HD-GCN as an extractor but also works on models with other structures like ST-GCN.
4.3.7 Detection result analysis and visualization
Based on the performance of the train and test sets of NTU-RGB+D 60, the energy distribution in the test set is bimodal. We suspect that some out-of-distribution classes are similar to in-distribution classes, making them difficult to distinguish. We list these ‘confusing’ classes in Table 10.
Table 10: Misclassification statistics
Wrong number 533 73 48 93 96 50
ID classes clapping giving object wipe face headache fan self pat on back
OOD classes hand waving, kicking something, rub two hands, neck pain, touch pocket
It displays the count of samples and their class names respectively from OOD samples that are misclassified into ID classes. Upon examining the class names, it is apparently that certain actions have subtle distinctions but still similar, such as ‘clapping’ and ‘hand waving’, or ‘pat on back’ and ‘neck pain’ as shown in Figure 7. How the model captures the little differences between these actions still needs some improvement.
(a) A10-ID: clapping
(b) A23-OOD: hand waving
(c) A53-ID: pat on back
(d) A47-OOD: neck pain
Figure 7: Example of correct classified ID samples and misclassified OOD samples.
@@ 5 Conclusion and future works
This paper introduces Skeleton-OOD, an end-to-end framework designed to address the challenge of recognizing out-of-distribution human action samples. The framework adopts a hierarchical graph composition method and utilizes GCN to extract the feature maps of human skeletons. Subsequently, a feature activation strategy is applied and we fuse the activated features with the original ones through a proposed feature fusion block. Experiments show that this step equips the model with the innate ability to identify out-of-distribution data while maintaining its capability for ID classification. Next, an energy score is computed for each test sample based on the logits output by the classifier. This helps determine whether it belongs to an OOD category. The model is trained using a designed energy-based loss function, designed to reduce the variance of ID data and differentiate between ID and OOD samples. Experimental results demonstrate that this approach effectively mitigates the overconfidence issue encountered by human action classification models when dealing with samples beyond the training category. Notably, since the model training process relies exclusively on sample information within the branch, without prior knowledge of OOD distribution characteristics, it exhibits broad practical applicability.
@@
Admittedly, the model still has some potential limitations. Firstly, since the model is trained on GCN, its generalization ability is constrained by the number of skeleton nodes across various datasets. In real-world scenarios, if the trained model is deployed to identify whether a test sample with a different number of nodes is out-of-distribution, its accuracy may be compromised. Secondly, the end-to-end model’s generalization capability is still bounded by the model parameters and the size of the dataset. The way the activation method is used during training causes some of the improvement in results to come from the training operation. One potential solution is to explore semi-supervised learning techniques to mitigate this issue. Moreover, the problem can be further addressed by introducing generative model and conducting training on datasets with limited ID samples. By leveraging existing category text and feature information, the model can infer the name of the unknown category for unfamiliar sample features. In summary, human action recognition tasks present diverse application scenarios, with numerous novel challenges awaiting exploration based on specific practical contexts.
Appendix A Appendix
A.1 Three different activation shaping strategies
In our paper, we try three different activation shaping strategies and compare the results respectively. Here are their implementation details in Table 11.
Table 11: Detailed activation shaping algorithms
Algorithm S1. Activation shaping algorithms
Input: Feature maps $F$;
      Pruning percentage $p$;
Output: Modified feature maps $F_{P}$, $F_{B}$, or $F_{S}$;
ASH-P: Activation Shaping with Pruning ASH-B: Activation Shaping with Binarizing ASH-S: Activation Shaping with Scaling
1: $t=p\times\operatorname{rank}(F)$ 1: $t=p\times\operatorname{rank}(F)$ 1: $t=p\times\operatorname{rank}(F)$
2: $F_{P}[\operatorname{idx}(\mathrm{F}<\mathrm{t})]=0$ 2: $s=\operatorname{sum}(F)$ 2: $s1=\operatorname{sum}(F)$
3: return $F_{P}$ 3: $F_{P}[\operatorname{idx}(\mathrm{F}<\mathrm{t})]=0$ 3: $F_{P}[\operatorname{idx}(\mathrm{F}<\mathrm{t})]=0$
4: $n=\operatorname{len}(\operatorname{idx}(F\neq 0))$ 4: $s2=\operatorname{sum}(F)$
5: $F_{B}[F\neq 0]=s/n$ 5: $F_{S}=F[\operatorname{idx}(F\neq 0)]\times\exp(s1/s2)$
6: return $F_{B}$ 6: return $F_{S}$